OpenStudy (anonymous):

Drug doses for dogs and cats are known to scale with their surface area S. When body mass W is measured in kilograms, then surface area S in square meters is given by S=K* W ^(2/3)/(100), where for dogs K=10.1 and for cats K-10.4. Further, when converting human drug doses of an average adult to pet drug doses, this formula is used: pet's drug dose =pet's S/(1.73) * human adult drug dose Calculate the drug dose (rounded to the nearest milligram) that you would give your dog or car of the indicated weight. 200 mg of aspirin and your cat weighs 4.6kg. I get 16.6 but the book says its 33 kg.

OpenStudy (jack1):

S=K* W ^(2/3)/(100) is this: \[S=K* W ^{(2/3)/(100)} \rightarrow or \rightarrow S=\frac {K* W ^{(2/3)}}{100}\] ??

OpenStudy (jack1):

assuming it's the 2nd one: S=K* W ^(2/3)/(100) pet's drug dose =pet's S/(1.73) * human adult drug dose S=K* W ^(2/3)/(100) K=10.4 W=4.6 kg so S = 0.287655 pet's drug dose =pet's S/(1.73) * human adult drug dose x = S/1.73 * 200mg ...again is this: \[x = \frac {S}{1.73 \times 200mg} \rightarrow or \rightarrow x = 200mg \times \frac S {1.73}\] ??? assuming the 2nd one: x = (200 x S)/1.73 x = 200 x 0.287655/ 1.73 x = 33.25 mg

OpenStudy (anonymous):

@Jack1 I see what I did wrong, what I did wrong was that for this part; S=K* W ^(2/3)/(100) K=10.4 W=4.6 kg so S = 0.287655, the part where it says to times it with 100 I times it with 200....oops. So from there I ended up getting the rest of the problem wrong. Thanks for clarifying.

OpenStudy (jack1):

ah, simple finger slip on the calc, all good hay, glad u solved it